IP A 2.8 kg block slides with a speed of 2.1 m/s on a frictionless horizontal surface until it encounters a spring. Part A If the block compresses the spring 5.6 cm before coning to rest, what is the force constant of the spring? Express your answer using two significant figures. N/m Submit Request Answer Part B What initial speed should the block have to compress the spring by 1.4 cm? Express your answer using two significant figures. UE m/s Submit
Answer #1

m = 2.8 Kg

v =2.1 m/s

part A) x = 0.056 m

let the force constant is k

using conservation of energy

0.50 * k * 0.056^2 = 0.50 * 2.8 * 2.1^2

solving for k

k = 3937 N/m

the spring constant is 3937 N/m

part B)

let the speed is v

using conservation of energy

0.50 * k * 0.014^2 = 0.50 * 2.8 * v^2

solving for v

v = 0.524 m/s

the initial speed is 0.524 m/s